OpenStudy (anonymous):

p, q , r , s, n and m are positive integers such that 4p = 5q = 6r = 12s = 8n = 9m which of the following pairs contains a number , which is not an integer? {p/m , m/q} {q/9 , rs/p}

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

its p/m and m/q u see both are coming in decimal if u solve them

OpenStudy (anonymous):

yes u r correct......how????

OpenStudy (unklerhaukus):

\[\frac pm =\frac 94\]\[\frac mq=\frac59\]\[\frac q9=\frac{1}{45}\]\[ \frac{ rs}{p}=\frac{4}{72}\]

OpenStudy (anonymous):

ok then...

OpenStudy (anonymous):

P/m=4/9 when u will olve u ll get anwer in decimal M/Q=9/5 here u also get in deciaml q/9=5/9 u see here answer is in deicaml but in its apir anwer is not in decimal rs/p=72/4 so this pair cant be the ans

OpenStudy (anonymous):

apir=*pair

OpenStudy (unklerhaukus):

i got q/9 wrong

OpenStudy (anonymous):

thanzz

OpenStudy (unklerhaukus):

im a bit confused with rs/p

OpenStudy (anonymous):

it happens :) thats y i always do qn first in my notebook :)